# Nighttime Person Re-Identification via Collaborative Enhancement Network with Multi-domain Learning

### v1

@@ 4.2 Comparison with State of the Arts
In this section, we compare CENet with six state-of-the-art ReID algorithms: IDE+ [43], PCB [23], ABD-Net [2] , AGW [31], BoT [19], and TransReID [9] on the Night600, $RGBNT201_{rgb}$ and Syn_Dark datasets. To ensure fair evaluation, we retrain all methods using nighttime data. Additionally, we compare the proposed method (CENet) with IDF [18], which is a state-of-the-art ReID method designed for night scenes. As shown in Table 2, CENet demonstrates a significant performance advantage over all competing algorithms in real nighttime datasets. When compared to IDF on both the Night600 and $RGBNT201_{rgb}$ datasets, CENet achieves impressive improvements of 4.1% in mAP and 7.8% in Rank1, as well as substantial gains of 17.3% in mAP and 19.0% in Rank1, respectively.
@@
Table 3: Comparison results of CENet with state-of-the-art multi-modality methods on RGBNT201 dataset
@@
Moreover, we evaluate CENet with state-of-the-art multi-modality ReID methods on the RGBNT201 dataset. As shown in Table 3, CENet only relies on RGB modality data and achieves performance that significantly surpasses that of the state-of-the-art multi-modal ReID algorithms. Specifically, CENet outperforms IEEE [26] by a significant margin, achieving improvements of $8.9\%$ in mAP and $10.3\%$ in Rank-1 metrics, which indicates that visible light images at night also contain huge potential.
4.3 Ablation Study
Table 4: Ablation studies of CENet on Night600 and Syn_Dark datasets
@@ 4.3 Ablation Study
CENet w/o MD PS 8.80 16.10 31.30 79.80 91.90 97.20
CENet w/o MD FD 8.90 17.20 31.50 76.20 90.00 96.20
CENet w/o MD FD PS 8.40 16.00 31.10 75.50 89.90 96.30
In this section, we evaluate the effectiveness of each component in CENet on two different datasets: a real nighttime ReID dataset (Night600) and a synthetic ReID dataset (Syn_Dark). We systematically isolate and evaluate the three components within CENet: parameter sharing (PS), feature distillation (FD), and multi-domain learning (MD). Specifically, we first exclude MD from CENet, labeling it as ”CENet w/o MD”. As we can observe from Table 4, the exclusion of MD leads to a notable decline in performance. Subsequently, we individually exclude PS and FD, labeling them as ”CENet w/o MD PS” and ”CENet w/o MD FD”. This also results in a degradation in model performance. Finally, we exclude all the components proposed in this paper, establishing it as the baseline method referred to as ”CENet w/o MD FD PS”. Further detailed results are presented in Table 4, which illustrates that each component plays an important role in nighttime person ReID.
To evaluate the effectiveness of the proposed lighting distillation (LD) loss, we also consider two commonly used feature distillation losses: mean squared error (MSE) loss and Kullback-Leibler (KL) loss for feature distillation. To this end, we construct three different variants, namely CENet-LIG w/o MD, CENet-MSE w/o MD, and CENet-KL w/o MD, which respectively indicate replacing LD with only lighting loss, MSE loss, and KL loss. From the results in Table 5, it can be observed that these variant losses do not bring performance improvement when applied to the real nighttime dataset Night600. This demonstrates the effectiveness of the proposed LD loss.
Table 5: Lighting distillation (LD) loss ablation studies on the Night600 dataset
@@
4.4 Visualization
We present the output results of the relighting branch based on the ”CENet w/o MD” and the ”CENet” in Figure 5. It can be observed that the ”CENet w/o MD” effectively enhances the brightness of low-light images, indicating that increasing the brightness of nighttime images contributes to improved nighttime ReID performance. However, the unsupervised learning approach in the relighting branch exhibits limitations in terms of color shift and noise smoothing, as illustrated in the second row of Figure 5. In contrast, CENet incorporates a supervised relighting learning process using synthetic data. CENet not only improves the image brightness but also enhances the color quality and smoothness of noise, leading to better performance. This highlights the importance of enhancing the color and noise quality of nighttime images for nighttime ReID performance.
Figure 5: Examples of relighting image under different models
Figure 6: Visualization analysis based on t-SNE technique
@@ 4.4 Visualization
In addition, we randomly selected 12 different samples from the test set and extracted their features using three models: baseline, CENet w/o MD, and CENet. Then, we visualized the distribution of the sample features using t-SNE technique [24]. Figure 6 (a) shows the representation distribution of these sample features by the baseline model, where different colored dots represent different identities. It can be observed that the sample features of different individuals in the region represented by the blue dashed line are difficult to distinguish. In contrast, CENet w/o MD demonstrates improved inter-class separation and intra-class cohesion, enhancing the overall feature distribution, as shown in Figure 6 (b). This indicates the effectiveness of the proposed network. Furthermore, with the introduction of multi-domain learning, the intra-class compactness is further enhanced, leading to improved discriminability of hard samples, as shown in the red dashed region in Figure 6 (c). This confirms the effectiveness of multi-domain learning.
5 Conclusion
@@
In this paper, we presents a novel solution for person ReID in nighttime scenes. By employing a parallel network architecture, we simultaneously perform image relighting and person ReID, mitigating the impact of relighting image quality on the ReID task. We design multi-level feature interaction based on parameter sharing and feature distillation to collaborative modeling two tasks. Furthermore, we propose that multi-domain learning helps enhance feature representation and improve model performance. In future, we plan to explore unsupervised learning schemes for nighttime ReID, allowing to leverage larger-scale unlabeled real nighttime data and further improving performance.

### v2

@@ IV-B Comparison with State-of-the-art Methods
In this section, we compare CENet with six state-of-the-art ReID algorithms, including IDE+ [55], PCB [56], ABD-Net [57] , AGW [59], BoT [58], and TransReID [4] on Night600, $RGBNT201_{rgb}$ and Syn_Dark datasets. To ensure fair evaluation, we retrain all methods using nighttime data. Additionally, we compare CENet with IDF [7], which is a state-of-the-art ReID method designed for night scenes. From the results in Table II, CENet shows significant performance advantages overall comparison algorithms on both real and synthetic nighttime person ReID datasets.
Evaluation on Night600. From the results in Table II, it can be seen that CENet achieves significant improvements of 4.1% on mAP and 7.8% on Rank-1 compared to IDF [7] on Night600 dataset. Notably, IDF is specifically designed for nighttime ReID. As a typical sequential combination of relighting and ReID networks, IDF relies on the quality of relighting images in the inference phase, and is thus limited in accuracy and efficiency. In contrast, CENet removes relighting subnet during the inference phase and directly performs ReID task on nighttime images. Therefore, these results demonstrate the significant superiority of the proposed CENet in accuracy and inference efficiency against other methods.
In addition, compared with the existing conventional ReID algorithms, CENet outperforms TransReID and AGW by 4.9%/9.0% and 7.2%/12.5% on mAP/Rank-1, respectively. In particular, the TransReID method, serving as the baseline for our approach, maintains a network structure identical to CENet during the inference stage. Therefore, the above experimental results strongly suggest the significant advantages of CENet in the night ReID task.
Evaluation on RGBNT201rgb. We compare our method with the advanced ReID algorithm on another real nighttime ReID dataset $RGBNT201_{rgb}$, and the results are shown in the Table II. Similarly, CENet still exhibits a clear performance advantage. Specifically, compared to the second best IDF algorithm on this dataset, CENet achieves a significant improvement of 17.3% and 19.0% on mAP and Rank-1, respectively. Compared with TransReID and AGW, CENet outperforms them by 19.2%/26.7% and 21.8%/36.3% on mAP/Rank-1 metrics respectively. Although all methods perform better on the $RGBNT201_{rgb}$ dataset than on the Night600 dataset, this improvement can be attributed to the smaller scale of the $RGBNT201_{rgb}$ test set. With only 30 identities, this dataset does not fully represent the complexity of nighttime conditions. In addition, by comparing the performance of two nighttime datasets of different sizes, we believe that further expanding the size of nighttime training data is the key to improving nighttime ReID performance.
Evaluation on Syn_Dark. In Table II, we show the experimental results of CENet and existing state of the art methods on the synthetic nighttime ReID dataset. It can be seen that compared with IDF and TransReID, CENet achieves 4.1%/3.0% and 3.7%/1.8% performance improvement on mAP and Rank-1, respectively, which indicates that the proposed method demonstrates effective modeling capability for synthetic domain data. Notably, IDF, which combines relighting network and ReID network in a sequential manner, performs lower on synthetic datasets than conventional ReID algorithms (TransReID), suggesting that sequential structures on synthetic domain data are difficult to avoid performance degradation due to unstable relighting results. On the other hand, it also reflects the difference between synthetic data and real data. Therefore, exploring low-light scene ReID only on the synthetic dataset is hard to reflect the performance of real nighttime scenarios. However, compared with daytime, the current nighttime ReID performance is still low, which indicates the limitations of existing night ReID methods.
Comparison with multimodal ReID. To deal with the nighttime ReID problem, one potential approach is to incorporate visual modalities that are less affected by low-light conditions, thus mitigating the problem of poor imaging quality in RGB low-light scenes. In this context, Zheng et al. [54] propose a multimodal dataset to explore night ReID, which attracts attention to the research of multimodal ReID algorithms. However, the popularization of multimodal monitoring devices and the registration of multimodal data require substantial costs, which limits the practical applications of multimodal ReID. To this end, we explore the potential of low-light RGB images. In Table III, we show the experimental results of CENet and existing multimodal ReID methods. It can be seen that CENet, which only uses RGB data for training and evaluation, performs much better than the algorithms that use multi-modal data for training and evaluation. Specifically, CENet achieves 8.9%/10.3% and 16.8%/18.5% performance advantages over IEEE and PFNet on mAP and Rank-1, respectively.
It is worth discussing that current multi-modal ReID methods primarily adopt CNN network to extract person features, which lack crucial global modeling capabilities. In contrast, CENet employs a powerful Transformer network for feature extraction and incorporates synthetic low-light data during training, significantly improving the model’s representation and generalization abilities. Consequently, CENet can achieve superior performance using only the RGB modality in low-light conditions. In addition, several Transformer-based multi-modal ReID algorithms [60, 61] published after this work achieve more superior performance, which still demonstrates the effectiveness of multi-modal information. In summary, this indicates that existing multimodal ReID algorithms have not fully exploited the potential of the RGB modality, and that ReID under nighttime RGB images has great potential.
@@
TABLE III: Comparison results of CENet with state-of-the-art multi-modality methods on RGBNT201 dataset
@@
Comparison with cross-modality ReID. To further validate the effectiveness of the proposed method, we compare it with current well-developed cross-modal ReID methods, including DEEN [21], DMA [28], HiCMD [19], and XIVReID [20]. We regard a copy of night600 as the NIR modality to make the cross-modal ReID algorithm executable.
TABLE IV: Comparison results of CENet with state-of-the-art cross-modality methods on Night600 dataset
Methods Pub. info mAP Rank-1 Rank-5 Rank-10
HiCMD [19] CVPR 2020 2.94 7.66 14.77 19.27
XIVReID [20] AAAI 2020 3.29 14.27 31.70 40.37
DEEN [21] CVPR 2023 12.84 21.56 36.97 45.69
DMA [54] TIFS 2024 12.75 20.31 35.63 43.83
CENet Ours 13.30 25.00 43.00 51.70
The results are shown in Table IV, and it can be seen that the current state-of-the-art cross-modal ReID algorithms, DMA and DEEN, both achieve notable performance in Night600. The reason can be attributed to the fact that these algorithms not only focus on the problem of modal differences, but also design the feature extraction network carefully. It indicates that improving the feature extraction capability of the network is one of the key factors to improve the performance of ReID at night. Therefore, the main reason for the poor performance of XIVReID and HiCMD in Night600 can be attributed to the fact that these two algorithms focus only on solving the cross-modal differences and neglect the careful design of the feature extraction network. Although advanced cross-modal methods achieve good performance, our method exceeds the mAP/Rank1 metrics by 0.55%/4.69% and 0.46%/3.44% compared to DMA and DEEN, respectively, a result that supports the superiority of the proposed method. In addition, we observe that the DMA and DEEN algorithms have better performance than existing studies of unimodal algorithms, such as TransReID, AGW, etc., which suggests that the existing studies of cross-modal ReID may be more important references for visible-based nighttime ReID.
TABLE V: Ablation studies of CENet on Night600 and Syn_Dark datasets
@@
CENet w/o MD MFI 8.40 16.00 31.10 75.50 89.90 96.30
CENet w/o MD MFIFD 8.80 16.10 31.30 79.80 91.90 97.20
CENet w/o MD MFIPS 8.90 17.20 31.50 76.20 90.00 96.20
IV-C Ablation Study
Effectiveness of components. We perform an in-depth evaluation of the key components of CENet on two different datasets, Night600 and Syn_Dark. The major components under investigation are Multilevel Feature Interaction (MFI) and Multi-Domain learning (MD). Initially, we remove MD from CENet and label this model as “CENet w/o MD”. As shown in Table LABEL:tab:tcte_ablation, this leads to significant performance drops on the Night600 dataset, with mAP and Rank-1 decreasing by 3.8% and 5.8% respectively. On the Syn_Dark dataset, mAP and Rank-1 declined by 1.9% and 0.7% respectively. It shows that MD plays a key role in the model’s learning in both the synthetic and real domains. Subsequently, we exclude both MD and MFI and label the model as “CENet w/o MD MFI”, which is our baseline method. From the table  LABEL:tab:tcte_ablation, it can be observed that the model performance in Night600 mAP decreased by 4.9%, Rank-1 by 9%, while mAP is down 3.7% and Rank-1 is down 1.8% on the Syn_Dark dataset. These results further demonstrate the effectiveness of the designed MFI for ReID networks.
In addition, to further analyze the importance of each component in the multistage feature interaction, we conduct experiments on two feature interaction strategies respectively. “CENet w/o MD MFIFD” stands for removing the multi-domain learning strategy and the feature distillation scheme of the multilevel feature interaction, to verify the effectiveness of parameter sharing. As shown in Table LABEL:tab:tcte_ablation, compared with the baseline method, “CENet w/o MD MFIFD” improves mAP and Rank-1 by 0.4%/0.1% and 4.3%/2% on Night600 and Syn_Dark datasets, respectively, verifying the effectiveness of the low-level feature interaction. Similarly, ”CENet w/o MD MFI PS” stands for removing the multi-domain learning strategy and the parameter-sharing strategy of the multilevel feature interaction. Compared with the baseline method, the mAP/Rank-1 of the ”CENet w/o MD MFI PS” on Night600 and Syn_Dark datasets are increased by 0.5%/ 1.2% and 0.7%/ 0.1% respectively, indicating the effectiveness of high-level feature interaction. In conclusion, the combination of the two interactions results in a more significant improvement, that is, an increase of 1.1%/3.2% and 1.8%/1.1% in the mAP/Rank-1 metrics on Night600 and Syn_Dark datasets, respectively.
TABLE VI: Feature distillation (FD) loss ablation studies on the Night600 dataset
@@ IV-C Ablation Study
Effectiveness of distillation loss. To evaluate the effectiveness of the proposed feature distillation (FD) loss, we also investigate two commonly employed distillation losses: mean squared error (MSE) loss and Kullback-Leibler (KL) loss for feature distillation. To this end, we construct three different variants, namely “CENet-BL w/o MD”, “CENet-MSE w/o MD”, and “CENet-KL w/o MD”, which respectively indicate replacing FD with only brightness loss, MSE loss, and KL loss. From the results in Table LABEL:tab:tcte_ablation2, it can be observed that these alternative losses do not bring performance improvement when applied to the real nighttime dataset Night600. This demonstrates the effectiveness of the proposed feature distillation loss.
TABLE VII: Ablation studies of multi-domain learning on Night600 and Syn_Dark datasets
Night600 Syn_Dark
mAP Rank-1 Rank-5 mAP Rank-1 Rank-5
Single-domain learning 9.50 19.20 33.50 77.30 91.00 96.60
Mix-domain learning 10.50 22.50 36.40 72.10 87.70 95.50
Multi-domain learning 13.30 25.00 43.00 79.20 91.70 97.00
Effectiveness of multi-domain learning. To evaluate the effectiveness of the proposed multi-domain learning, we also investigate two commonly employed learning scheme: single-domain learning and mix-domain learning for network training. Specifically, we design a single-domain learning scheme that trains exclusively on the real-domain dataset, and a mixed-domain learning scheme that trains on a combined dataset of real and synthetic domains. The multi-domain learning approach exhibits clear advantages, as reflected in Table LABEL:tab:tcte_ablation_MDL. It surpasses the single-domain learning approach by 3.8% and 4.8% in mAP and Rank-1 metrics on the real domain test set. Moreover, when compared to the mixed-domain learning approach, multi-domain learning strategy demonstrates superior performance with a 2.8% and 2.5% increase in mAP and Rank-1 metrics, respectively. These results further validate the effectiveness and superiority of the multi-domain learning strategy.
TABLE VIII: Ablation study of the importance of parallel architecture
Traning dataset Testing dataset Improvement
Night600 Syn_dark Rank-1 mAP
Baseline 16.00 8.40
w/ Data 18.10 8.90 2.1/0.5
w/ Parallel 19.20 9.50 3.2/1.1
w/ Data Parallel 22.50 10.50 4.4/1.6
w/ Data Parallel MD 25.00 13.30 2.5/2.8
w/ Data Sequential 20.30 10.00 2.2/1.1
w/ Data Sequential MD 20.10 10.50 -0.2/0.5
Importance and necessity of parallel architecture. Since extra data and multi-domain learning are integrated into the parallel architecture, it is difficult to clearly demonstrate their benefits. Therefore, we conduct a series of experiments to illustrate the core impact of the parallel architecture. Specifically, we compare the parallel architecture (Parallel), extra data (Data), and multi-domain learning (MD) in two perspectives. First, each component is introduced incrementally in the baseline approach. As shown in Table VIII, the “w/ Parallel” and “w/ Data Parallel” show the parallel architecture improves Rank-1/mAP by 3.2%/1.1% and 4.4%/1.6% over the baseline method TransReID before and after the introduction of extra data, respectively. The “w/ Data” indicates that introducing extra data in TransReID brings 2.1%/0.5% improvement over TransReID. The “w/ Data Parallel MD” indicates that improves 2.5%/2.8% on Rank-1/mAP in the parallel architecture. Second, we introduce extra data and multi-domain learning into the sequential framework (Sequential) to further demonstrate the advantages of parallel framework. As shown in the last two rows of Table VIII, although the performance gains from extra data and multi-domain learning in the sequential framework are effective, it is significantly less than the parallel framework, which emphasizes the importance and necessity of the parallel framework.
Efficiency analysis of training and testing phases. We also provide efficiency analysis of training and testing phases between CENet and its variants. In training phase, we compare the training time of CENet with and without extra data in Table IX. It can be seen that introducing extra datasets significantly increases the training time. Specifically, training with only Night600 takes about 34 hours, while adding synthetic data extends it to 87 hours. Similarly, training with only RGBT201 takes about 3 hours, but with extra synthetic data, it extends to 72 hours. Notably, due to training on a public server, the training time is also related to the number of programs running on the server.
TABLE IX: Ablation study of model training time before and after introduction of extra data
Training time CENet CENet w/ Syn_dark Platform
Night600 34 hours 87 hours RTX3090*1
RGBNT201 3 hours 72 hours
In testing phase, we present the inference time comparison between with and without the relighting branch in Table X. It can be seen that remove relighting branch saves 14 seconds and 33 seconds in night600 and Syn_drak test datasets, respectively. Notably, the Syn_dark test set is larger than the Night600 test set, leading to greater time savings for larger test sets.
TABLE X: Ablation Study of time savings before and after removing the relighting branch
Inference time Night600 Syn_dark Platform
w/ relighting 85 seconds 181 seconds RTX3090*1
w/o relighting 71 seconds 148 seconds
IV-D Visualization
Figure 6: The t-SNE [64] visualization of features on Night600 dataset.
@@ IV-D Visualization
Visualization of feature distribution. To visualize the distribution of sample features, we randomly select 12 different samples from the test set and extract their features using three models: baseline, CENet w/o MD, and CENet. Then, we employed the t-SNE technique [64] to project the high-dimensional features into a two-dimensional space. Figure 6 (a) illustrates the representation distribution of these sample features by the baseline model, where different colored dots represent distinct identities. It can be observed that the sample features of different individuals within the blue dashed line region are difficult to distinguish. In contrast, CENet w/o MD demonstrates improved inter-class separation and intra-class cohesion, enhancing the overall feature distribution, as depicted in Figure 6 (b). This demonstrates the effectiveness of the proposed network. Furthermore, the introduction of multi-domain learning in CENet further enhances intra-class compactness, leading to improved discriminability of hard samples, as highlighted by the red dashed region in Figure 6 (c). This confirms the effectiveness of multi-domain learning in enhancing the feature representation for challenging samples.
Figure 7: Visualization of relighting images from three models, including existing advanced relighting method Zero-DCE [9], CENet without multi-domain learning and CENet with multi-domain learning
Visualization of relighting images. We present the original nighttime images, the images relighted by the advanced relighting algorithm Zero-DCE [9], and output results of the relighting branch based on the “CENet w/o MD” and the “CENet” in Figure 7. It can be seen that the enhancement quality of “W/ Zero-DCE” is relatively low, especially in terms of brightness and color. It also can be observed that the “CENet w/o MD” effectively enhances the brightness of low-light images, indicating the significance of increasing image brightness for improved nighttime ReID performance. However, the unsupervised learning approach in the relighting branch limits color shifting and noise smoothing, as illustrated in the second row of Figure 7. In contrast, CENet incorporates a supervised relighting learning process using synthetic data. CENet not only improves the image brightness but also enhances the color quality and smoothness of noise, leading to better performance. This highlights the importance of enhancing the color and noise quality of nighttime images for nighttime ReID performance. However, due to the influence of noise, the overall image quality still has more room for improvement.
Figure 8: Visualization of sample ReID results from three models.
Visualization of sample ReID results. We also visualize the recognition results of a selection of representative samples to verify the effectiveness of the proposed method. In Figure 8, we present the recognition results of two examples under three models: “Baseline”, “CENet w/o MD” and “CENet”, respectively. It can be seen that the “Baseline” model mismatches a large number of very low light samples due to the difficulty in distinguishing them. However, in the results of “CENet w/o MD” model, the matching accuracy is effectively improved, even for some low-light samples. In “CENet”, the matching results are further improved, which shows the effectiveness of the proposed method. In addition, it can be observed in the demonstrated examples that the same person has obvious color shift problems in the night scene, which brings certain challenges to the conventional ReID method. However, our method can effectively avoid its influence, which indicates the necessity of studying the nighttime ReID algorithm.
V Conclusion
@@
We present a novel solution to address the challenge of person ReID in nighttime scenes. Our method leverages a parallel network architecture to simultaneously perform image relighting and person ReID, effectively avoiding the impact of relighting image quality on the ReID task. To achieve effective enhancement of ReID features from the relighting task, we design multilevel feature interaction based on parameter-sharing and feature distillation, which enhance both low-level and high-level features in ReID. In addition, we design a multi-domain learning scheme that learns more robust feature representations from data across different domains, significantly improving the model’s performance in real night scenes. In future work, we aim to explore unsupervised learning schemes for nighttime ReID, which will be able to leverage larger-scale unlabeled real nighttime data, thereby further improving the performance of nighttime ReID.